OpenStudy (katielong):

Integrate: { -3/x +6/x^2 + 3/(x+3) dx (With { dx being the notation for integration, I can't find the actual sign, and with 2 over 1 as the x substitutes)

OpenStudy (anonymous):

So is that \[\int_1^2\left(-\frac{3}{x}+\frac{6}{x^2}+\frac{3}{x+3}\right)\,dx~~?\]

OpenStudy (katielong):

Yes

OpenStudy (anonymous):

For the first and third terms, you would use the logarithm, since \[\int \frac{dx}{x}=\ln|x|+C\] The second integral uses the power rule: For \(n\not=-1\), \[\int x^n\,dx=\frac{x^{n+1}}{n+1}+C\]

OpenStudy (katielong):

Okay so it should be: [-3lnx - 6x^-1 + 3ln(x+3) ] ?

OpenStudy (anonymous):

That's right. Now you need to evaluate the endpoints as you would with any definite integral. If \(F(x)=-3\ln|x|-\dfrac{6}{x}+3\ln|x+3|+C\), find \(F(2)-F(1)\).

OpenStudy (katielong):

Okay, thank you. = (-3ln2-3+3ln5)-(-3ln1-6+3ln4) How would this be simplified? It supposedly goes to 3-2ln(8/5) ?

OpenStudy (anonymous):

Close, that 2 should be a 3.